I have a problem with PS100+LS336 combination. After the installation was moved to another building the sensor on sample and that on the radiation shield show different readings. At room temperature the first shows 294 K (probably correct) and the second 318-321 K. Can somebody say what can be the reason and how bring it back to work?
Post by Lake Shore Jeff M on Mar 2, 2021 8:49:49 GMT -5
I will start by explaining how the 336 determines the temperature of a PT-100 sensor. To start it sources 1mA of excitation current on the I+ & I- leads and then reads the voltage across the V+ & V- leads. Then it determines the resistance by dividing the voltage by the current followed by a look up in the curve table. using the curve table, it locates the breakpoint values that are above and below the measured resistance and then does a straight line interpolation between those points to determine the temperature.
Using the information you provided, at 294K, the PT-100 sensor should be measuring around 108 ohms. If it is showing 318K, then the 336 is seeing a resistance around 116 - 120 ohms.
The 8 to 12 ohm difference is possibly due to a broken or shorted lead in the sensor wiring so let me know the results of the measurements. It may be best to send the results to email@example.com and provide the information regarding the system the sensor is installed in so we can provide additional troubleshooting support.
Post by Lake Shore Jeff M on Mar 4, 2021 8:31:58 GMT -5
Tei, the operation of the 336 is the dame for the Diode sensor as it is for the Platinum sensor. Can you tell me the serial number of your PS-100 probe station so I can put together a detailed troubleshooting procedure. Please email the information to firstname.lastname@example.org and i will respond accordingly.
Dear Jeff, Thanks for your efforts to help me. I have solved the problem. Just for the colleagues who might experience the same, I will give a short explanation. On PS100 probestation the connector to the thermoprobes (diodes) is located at the bottom of the setup, with very unhandy access. Probably the person who was connecting the 336 back didn't take enough care and as a result one of male lags ("H") was leaning out and not connected. This was the V- contact from the radiation shield. As a result the four probe scheme was not working, there was a positive shift in temperature and instability. First, I have just checked all cables (no problems detected) and reached the connector. I took proto by a cellphone and saw the leaning out contact. It was not so difficult to put the leg back and connect everything properly. Now everything seems to work properly. Best wishes, Tei